OpenStudy (anonymous):

Explain, in complete sentences, the effects of multiplying a radical or a complex number by its conjugate. Give an example of each.

OpenStudy (anonymous):

The reason we do rationalization or multiply by conjugate is because we don't want a radial or complex number in the denominator. It makes the result somewhat easier to read mathematically.

OpenStudy (anonymous):

Suppose you have 2 / i (*i is imaginary number which is sqrt(-1)) you multiply both numerator and denominator by i because i^2 = -1 so you have 2i/-1 = -2i

OpenStudy (anonymous):

2/sqrt(2) again you multiply by sqrt(2) for both numerator and denominator to get 2sqrt(2) / 2

OpenStudy (anonymous):

If you have something like [(i + 3)] or [sqrt(2) + 5] in the denominator of an algebraic expression, The conjugate will be [(i - 3)[ and [sqrt(2) - 5] respectively its just the same expression with opposite sign.

OpenStudy (anonymous):

You will multiply the conjugate both by numerator and denominator as well

OpenStudy (anonymous):

\[[2 / (i-3)]x[(i+3)/(i+3); = 2(i + 3)/[i ^{2}-9] = 2(i + 3)/[-10] = -(i+3)/5\]

OpenStudy (anonymous):

multiplying a number \[a+bi\] by its complex conjugate \[a-bi\] gives \[a^2+b^2\] a real number for example \[(4+3i)(4-3i)=4^2+3^2=25\]

OpenStudy (anonymous):

multiplying a radical \[a+\sqrt{b}\] by its conjugate \[a-\sqrt{b}\] gives \[a^2-b\] for example \[(4+\sqrt{3})(4-\sqrt{3})=4^2-3=13\]